New ZOOM room:

News:

Please fill the forms with your opinions on L2 candidates We need to fill the INFN consuntivi. Covid: had a meeting called by the INFN director. the law says that the ppl should stay at home as much as possible. Marta has asked to have multiple ppl in a room unless necessary. Legnaro forbids missions from PD unless discussed. Presence of technicials will be limited in the next 2 weeks. Need to discuss time shifts if we want to access the lab.

Framework & production

All the virtual machines have to be turned off. If you do not need them again please kill them By the end of March condor will replace LSF.

Analysis

B physics

PLB published!

Heavy composite neutrinos

Matteo: March 31st is the date for the dissertation. New recipe for high pt mu and fat jets applied and approved in the POG

VBS

Things with FNAL going on. Might need to move to ultralegacy.

WH_SS

Ugo: SY is finilazing his thesis, will freeze on March 31st. Ugo and Alberto sent comments.

H/Z->JPsi gamma

Rocco: signal generation approved. Cards approved means the gridpacks can be used to regenerate the signal. Alberto: Moving to ultralegacy. Data available but no MC available yet.

Higgs

Just got the datacards. We split excusively the 2 phasespaces w/o overlaps. Will start testing the combinations by the end off the week. Very high Pt where 2 jets are merge in a fat jet (pt_H ~200 GeV min), then low pt with 2 separate charm-tagged jets (can reach pt_h up to 400-500 GeV). The split is now around 250 GeV. If everything goes well we can ask for preapproval.

EPR

Detector

Trigger

Jacopo: advanced with the studies. Ready to report during the next DT trigger and software meeting, next Thursday at 10am. Marco: In the dept there is the plan to place a mu telescope. The idea is to use a few of the minichambers produced in Legnaro. Ugo: there is a first version of the project for the telescope facility. Mariotti, Collazuol.

Timing Layer

TEC: waiting for the delivery of small connectors, once soldered will be able to check the PCBs. Some modifications to be done to the cooling circuit. Daniele working on the current system, but will need to move the rack close to the fridge.

DT

Fabrizio: Franco should get back the fibre from the Chemistry dept. the fibre stayed 10 days downstream of MB3. Sandro: on the upgrade side things going ahead slowly.

Tracker/tracking

nothing to report.

A.O.B.

-- RobertoRossin - 2021-03-16

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2021-03-16 - RobertoRossin
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback